# FedALT: Federated Fine-Tuning through Adaptive Local Training with Rest-of-World LoRA

### v2

@@ F.3 Convergence Theorem
$\frac{\lambda}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (23)
@@
$L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\leq\frac{\mu}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$ (24)
@@
$\frac{\lambda}{2}\|Z_{k}^{t,s+1}-Z_{k}^{*}\|^{2}$ $\leq L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (25)
$\leq(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (26)
@@
$\|Z_{k}^{t+1}-Z_{k}^{t}\|$ $\leq\|Z_{k}^{t+1}-Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|$ (32)
$+\|Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-Z_{k}^{*}(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})\|$ (33)
$+\|Z_{k}^{*}(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})-Z_{k}^{t}\|$ (34)
@@
$\|(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})\|$ $=\frac{1}{K-1}\|\sum_{m\neq k}(Z_{m}^{t}-Z_{m}^{t-1})\|$ (36)
@@
Appendix G Limitations of FedALT in Traditional FL

### v3

@@ Appendix G Convergence Analysis
$\frac{\lambda}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$
$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (21)
@@
$L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\leq\frac{\mu}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$ (22)
@@
$\frac{\lambda}{2}\|Z_{k}^{t,s+1}-Z_{k}^{*}\|^{2}$
$\leq L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (23)
$\leq(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\quad-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (24)
@@
$\|Z_{k}^{t+1}-Z_{k}^{t}\|$
$\leq\|Z_{k}^{t+1}-Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|$
$\quad+\|Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-Z_{k}^{*}(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})\|$
$\quad+\|Z_{k}^{*}(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})-Z_{k}^{t}\|$ (30)
@@
$\|(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})\|$
$=\frac{1}{K-1}\|\sum_{m\neq k}(Z_{m}^{t}-Z_{m}^{t-1})\|$ (32)
@@
Appendix H Limitations of FedALT in Traditional FL